# Agentic Large Language Models, a survey

### v2

@@
Abstract
There is great interest in agentic LLMs, large language models that act as agents. We review the growing body of work in this area and provide a research agenda.
Agentic LLMs are LLMs that (1) reason, (2) act, and (3) interact. We organize the literature according to these three categories. The research in the first category focuses on reasoning, reflection, and retrieval, aiming to improve decision making; the second category focuses on action models, robots, and tools, aiming for agents that act as useful assistants; the third category focuses on multi-agent systems, aiming for collaborative task solving and simulating interaction to study emergent social behavior. We find that works mutually benefit from results in other categories: retrieval enables tool use, reflection improves multi-agent collaboration, and reasoning benefits all categories.
We discuss applications of agentic LLMs and provide an agenda for further research. Important applications are in medical diagnosis, logistics and financial market analysis. Meanwhile, self-reflective agents playing roles and interacting with one another augment the process of scientific research itself. Further, agentic LLMs may provide a solution for the problem of LLMs running out of training data: inference-time behavior generates new training states, such that LLMs can keep learning without needing ever larger datasets. We note that there is risk associated with LLM assistants taking action in the real world, while agentic LLMs are also likely to benefit society.
Website: https://askeplaat.github.io/agentic-llm-survey-site/
@@ Abstract
1 Introduction
The strength of the language abilities of LLMs has taken the world by storm. Recent work has extended their abilities with reasoning, information retrieval, and interaction tools. As a result, LLMs are now increasingly able to act as agents in the world (Shen, 2024; Qin et al., 2023). This ability has increased the relevance of LLMs to society and science. Agentic LLMs are being used to assist in medicine, logistics, finance, and other applications. Their ability to self-reflect, play roles, and interact enables new types of research, including large-scale social science simulations. We survey the growing body of literature on agentic LLMs, which we define as large language models that (1) reason, (2) act, and (3) interact. We organize this paper accordingly.
Agentic LLMs are also relevant in the acquisition of new training data for artificial intelligence (AI). Traditionally, LLMs have been trained on large datasets. Recently, however, it is getting harder to scale and improve datasets further, and training performance is reportedly plateauing, at high energy cost (Sutskever, 2024). By interacting with the world, agents generate new empirical data (see Figure 1). This data can be used for additional training (pretraining or finetuning) or to enhance performance at inference time, provided there is adequate grounding through human or automated validation and filtering (Subramaniam et al., 2025). An example of how LLMs can be trained based on their own actions, are Vision-Language-Action models, that update weights according to robotic action-feedback sequences (Black et al., 2024; Chiang et al., 2024; Yang et al., 2025). Thus, in addition to enabling useful applications, a second driver of interest in agentic LLMs is the opportunity to generate more training data.111Cognitive science teaches us that humans become more intelligent through interaction with the world and with other humans (we learn new behaviors and ideas from others). Societies of agents allow agentic LLMs to become more intelligent through interaction—as we will see in Section 4.
Agentic LLMs depend on progress in natural language processing, reasoning models, tool integration, reinforcement learning, agent-based modeling, and social science. At the confluence of these fields many exciting publications have emerged.
This paper has the following contributions:
@@ 1 Introduction
1.1 Agentic LLMs: Reasoning–Acting–Interacting
@@
Agency is a central concept in artificial intelligence (Russell and Norvig, 2016). Agency is about identity and control, and about the capability to act on one’s goals or will (Epstein and Axtell, 1996; Gilbert, 2019; Barker and Jane, 2016). Agents are endowed with decision-making capabilities, they communicate, sense changes in the environment, and act upon those changes (Wooldridge, 1999), see also Figure 2. Agents have been studied for a long time, and occur in many fields of AI. From the definition of agents interacting with the environment, different approaches focus on specific aspects of agents and agent behavior. In symbolic reasoning (Harman, 1984) and game theory (Von Neumann and Morgenstern, 2007; Owen, 2013), the topic of study is decision making by rational agents. The field of multi-agent systems studies intelligent systems that emerge from the interaction with different agents, human and/or artificial (Ferber and Weiss, 1999; Steels, 2003). In machine learning, the field of reinforcement learning studies how an agent can learn from interacting with an environment (Sutton and Barto, 2018). In this context, agents are systems that would adapt their policy if their actions influenced the world in a different way (Kenton et al., 2023). In autonomous systems and robotics, agents act in order to achieve a goal (Liu and Wu, 2018). Connectionism studies the emergence of intelligent behavior by embodied agents (Brooks, 1990; Medler, 1998). Evolutionary algorithms (Bäck, 1996) study nature-inspired computation by agents, of which the ant-colony optimization algorithm (Dorigo et al., 2007) is a well known example.
@@
The categories reasoning–acting–interacting build upon each other: the technology that has been developed by the reasoning approaches (category 1) is used for increasingly intelligent acting by assistants. The interactive abilities of the assistants (category 2) enable social simulation experiments. The outcome of assistant actions (category 2) and of these social experiments (category 3) can be used for data augmentation (category 1), to finetune LLMs (which can improve the accuracy of reasoning LLMs, etc.). This virtuous circle is depicted in Figure 1, and attracts interest from LLM researchers to agentic LLM methods (Sutskever, 2024; Guo et al., 2025; Du et al., 2025). The categories also correspond to fields in artificial intelligence that have a long research tradition across symbolic AI, robotics/autonomous systems, and connectionism/multi-agent modeling, respectively. Agentic LLMs are thus both a recent development and build on decades of research. This is reflected in our discussion below.
1.2 Literature Selection
@@
Related surveys on agents and LLMs are starting to appear. Li (2024) reviews retrieval and tool use in agentic LLMs. Wang et al. (2024b) focuses on autonomy and agent construction. Gao et al. (2024) also provide an extensive overview, and focus on multi-agent modeling and simulation. Xi et al. (2023) again focus on the construction of interactive agents, using a more explanatory anthropomorphic approach of perception, brain, and action. An extensive general survey of LLMs is Zhao et al. (2023), a slightly smaller one is (Minaee et al., 2024), an earlier survey is (Min et al., 2023). Yin et al. (2023) reviews works on multimodal LLMs.
@@ 1.3 LLM Training Pipeline
Originally, language models used recurrent architectures such as LSTMs (Hochreiter, 1997) to embed semantic relations between token structures, allowing limited connections between tokens. The transformer architecture is an effective implementation of the attention mechanism (Vaswani et al., 2017), allowing efficient random connections between tokens, improving performance greatly. Encoder transformer models, such as the BERT family (Devlin et al., 2018), learn embeddings that are suitable for text understanding and classification. Decoder transformer models, such as the GPT family (Brown et al., 2020), are trained by masking for text completion and instruction following, and are suitable for text generation.
@@
LLMs are trained on large datasets (Radford et al., 2019; Wei et al., 2022a). Performance on benchmarks testing formal linguistic competence is high (Warstadt et al., 2019) and so is accuracy on functional competence or natural language understanding tasks (GLUE, SQUAD, Xsum) (Wang et al., 2018, 2019; Rajpurkar et al., 2016; Narayan et al., 2018), translation (Kocmi et al., 2022; Papineni et al., 2002; Sennrich et al., 2015), question answering (Roberts et al., 2020). Even creative domains such as poetry composition LLMs have made progress by some standards (Zhang and Eger, 2024).
@@
Popular LLMs are OpenAI’s ChatGPT series (Achiam et al., 2023; Ouyang et al., 2022), Meta’s LLaMa family (Touvron et al., 2023), Anthropic’s Claude family (Anthropic, 2024), Google’s PaLM (Chowdhery et al., 2023) and Gemini (Anil et al., 2023), and the open-source models BLOOM (Le Scao et al., 2023), Pythia (Biderman et al., 2023), OLMo Groeneveld et al. (2024), and many others.
@@
LLMs are constructed using an elaborate pipeline with different training phases (Radford et al., 2019; Minaee et al., 2024). We will briefly describe the phases.
1. Acquire a large, general, unlabeled, text corpus (Brown et al., 2020).
2. Pretrain a transformer model on the corpus. This step yields a generalist natural language model. The pretraining is done using a self-supervised transformer approach (Vaswani et al., 2017) on the unlabeled dataset (text corpus).
3. Finetune the general model to a specific (narrow) task using a supervised approach on a labeled dataset consisting of prompts and answers (supervised finetuning, SFT) (Wei et al., 2022a; Minaee et al., 2024). This task can be, for example, translation from one language to another, or questions answering on a certain domain, such as medicine.
4. Instruction tune for improved instruction following. This is a form of supervised finetuning (Ouyang et al., 2022) to improve the ability to answer prompts.
5. Align the finetuned model with user expectations (preference alignment). The goal of this step is to improve the model to give socially acceptable answers such as prevention of hate speech. Popular methods are reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) or direct preference optimization (DPO) (Rafailov et al., 2024).
6. Optimize training to improve cost-effectiveness, for example with low-rank optimization (LoRa) (Hu et al., 2021), mixed precision training (Micikevicius et al., 2017), or knowledge distillation (Xu et al., 2024b; Gu et al., 2023).
7. Infer using natural language prompts (instructions). This stage, inference, is the stage where, finally, we can use the fruits of our training efforts. Prompting is the preferred way of using LLMs. In LLMs whose size is beyond hundreds of billions parameters a new learning method emerges: in-context learning (Brown et al., 2020; Wei et al., 2022a). This method provides a prompt that contains a small number of examples together with an instruction; it is a form of few-shot learning. However, no parameters of the model are changed by in-context learning, in-context learning takes place at inference time (Dong et al., 2022; Brown et al., 2020).
1.4 The Need for Agentic LLMs
@@
1. Prompt engineering Originally LLMs were trained as straight decoders, to be used with instruction prompts. The prompts contain context and instructions, and the model replies. The user interacts directly with the model, and writes the prompts themselves. LLMs turned out to be quite sensitive to small differences in the prompt formulation. When an answer is not satisfactory, the user has to remember the history of the interaction, and has to improve the prompt. This is known as prompt engineering. With basic LLMs, prompt-improvement is a tedious, manual, task.
2. Hallucination When LLMs provide answers that look good, but are factually incorrect, they are said to hallucinate. Hallucination is a major problem of LLMs. It is caused, in part, by a lack of grounding. Since LLMs are aligned to human preferences during fine-tuning, they often provide answers that look good by these standards while not adhering to other criteria, such as factuality. Various methods have been developed to mitigate hallucination, such as detecting uncertainty through self-reflection on its own answers, and with mechanistic interpretability methods. We will review papers that discuss these method in this survey.
3. Reasoning Another well-reported challenge for LLMs is (mathematical) reasoning (Plaat et al., 2024). LLMs used to be quite bad at solving math word problems (such as: “Annie has a one pie that she cuts into twelve pieces. She eats one third of the pieces. How many pieces does she have left?”). Reasoning challenges have given rise to step-by-step problem solving methods, such as reported by Wei et al. (2022b), both implicit, and with explicit (neurosymbolic) prompt optimization methods (Yao et al., 2024). This too we discuss in the next section.
4. Training Data LLMs are as smart as the data allows that was available at training time. When datasets no longer improve, pretraining and finetuning can no longer improve language models, and other learning methods are needed (Sutskever, 2024). Any event that happened after training, or any information available in special databases, are not in the model (Lewis et al., 2020).
@@ 1.5 Taxonomy
[…]
As we noted before, the three categories in our taxonomy come from three different backgrounds. To be intelligent, LLMs are enhanced with reasoning, combining deep learning with the symbolic AI tradition (Yu et al., 2024a; Li et al., 2025). To be active, LLMs are enhanced with tools that can act in the world (including robots, that plan to move in the world). To be social, LLMs are placed in interactive settings with other agents. They rely partially on capacities already present in traditional LLMs, such as basic theory of mind abilities and understanding of game theory and social dilemmas. Agentic LLMs learn to interact better by adapting their intelligence.
@@
Earlier progress in reasoning LLMs and retrieval augmentation has enabled much of the current developments in agentic LLMs (category 1). In this category, the aim is to address challenges in solving math word problems, and in providing up to date answers to queries. The contributions to intelligent LLMs came out of the need to improve reasoning of basic LLMs (sub-category a). In reasoning LLMs, methods from planning and search are used to let the LLM follow a step-by-step reasoning path. More elaborate search algorithms allow automated prompt-improvement and self reflection (sub-category b). Finally, certain questions can only be answered by inference-time data retrieval (sub-category c). The focus is on the individual improvement of the intelligent LLM agent.
@@
Category 3 is about interaction in multi-agent simulations. Here, in sub-category a, we first study basic social capabilities of LLMs on which interactions can build. Second, in sub-category b, we study how LLM agents can work together using simulations where they are assigned specific roles. Third, in sub-category c, we study emergence of collective phenomena in open-ended interactions, such as social coordination via conventions and norms. Here the focus is on the emergent level of the agent society. Multi-agent simulation with LLMs is becoming an active field for studying questions from the social sciences that previous generations of agent-based models were unable to address.
@@
| Chain of Thought (Wei et al., 2022b) | Step-by-step prompts | Math word (Cobbe et al., 2021) | Benchmark |
| Zero-shot CoT (Kojima et al., 2022) | ”Let’s think step-by-step” | Math word (Cobbe et al., 2021) | Benchmark |
| Self Consistency (Wang et al., 2022) | Ensemble | Math word (Cobbe et al., 2021) | Benchmark |
| Tree of Thoughts (Yao et al., 2024) | depth-first-search prompts | Game of 24 | Benchmark |
| Implicit Planning (Schultz et al., 2024) | Train SoS (Gandhi et al., 2024) | Chess, Hex | Benchmark |
| Progress Hint Prompt (Zheng et al., 2023) | Self Reflection | Math word (Cobbe et al., 2021) | Benchmark |
| Self Refine (Madaan et al., 2023) | Self Reflection | Dialogue Response | Benchmark |
| ReAct (Yao et al., 2022) | Reinforcement Learning | Decision Making | Benchmark |
| Reflexion (Shinn et al., 2024) | Self Reflection/Reinf Learning | Decision Making | Benchmark |
| Self Discover (Zhou et al., 2024a) | Self Reflection | Big Bench H (Suzgun et al., 2022) | Benchmark |
| Buffer of Thoughts (Yang et al., 2024c) | Self Reflection | Math word (Cobbe et al., 2021) | Benchmark |
| Memory Coordination (Zhang et al., 2023b) | Self Reflection | LLM Personalization | Benchmark |
| Adaptive Retrieval (Asai et al., 2023) | Adaptive Retrieval | Question Answering | Benchmark |
| Retrieval Augmentation (Lewis et al., 2020) | Retrieval Augmentation | Question Answering | Benchmark |
| MathPrompter (Imani et al., 2023) | Python Interpreter | Math problems | Benchmark |
| Program Aided Lang (Gao et al., 2023b) | Python Interpreter | Math word problems | Benchmark |
| Self Debugging (Chen et al., 2023d) | Debugger | Code generation | Benchmark |
| FunSearch (Romera-Paredes et al., 2024) | Genetic Algorithm | Algorithm Generation | Benchmark |
| Planning Language (Bohnet et al., 2024) | Planner/PDDL | Blocksworld | Benchmark |
| Self taught Reasoner (Zelikman et al., 2022) | Reason augm finetuning | Math Word | Benchmark |
| DeepSeek R1 (Guo et al., 2025) | Intrinsic Reasoning | Math Word | Benchmark |
@@
| WorldGPT (Ge et al., 2024) | Multimodal | World Model | WorldNet real-life scenarios |
@@
| WorldCoder (Tang et al., 2024) | Code model | World Code Model | Sokoban, MiniGrid, AlfWorld |
@@
| Task-planning (Guan et al., 2023) | PDDL World model | Task finetuning | AlfWorld |
| CLIP (Radford et al., 2021) | Multimodal | Vision Language | Benchmark |
@@
| Embodied BERT (Suglia et al., 2021) | Multimodal | Vision Language | ALFRED (Shridhar et al., 2020) |
@@
| E2WM (Xiang et al., 2024) | Embodied World Model | MCTS + World Model | Question Answering |
| RT-2 (Brohan et al., 2023) | Vision Language Action | VLA | Embodied reasoning tasks |
| LM-nav (Shah et al., 2023) | Action traces | VLA | Topological navigation |
| Mobility VLA (Chiang et al., 2024) | long context demonstration | VLA | Navigation MINT |
| (Black et al., 2024) | Flow Matching | VLA | Laundry folding, Table cleaning |
| Say Can (Ahn et al., 2022) | Grounded Actions | Value function for LLM | Manipulation, Kitchen |
@@
| Inner Monologue (Huang et al., 2022) | Grounded Actions | Affordance in prompt | Manipulation, Kitchen |
@@
| Lang Guided Expl (Dorbala et al., 2023) | Generic Class Labels | Vision/Language | L-ZSON |
| Automatic Tool Chain (Shi et al., 2024a) | grounded reasoning | Tool behavior | ToolFlow |

### v3

@@
Abstract.
Background: There is great interest in agentic LLMs, large language models that act as agents.
Objectives: We review the growing body of work in this area and provide a research agenda.
Methods: Agentic LLMs are LLMs that (1) reason, (2) act, and (3) interact. We organize the literature according to these three categories.
Results: The research in the first category focuses on reasoning, reflection, and retrieval, aiming to improve decision making; the second category focuses on action models, robots, and tools, aiming for agents that act as useful assistants; the third category focuses on multi-agent systems, aiming for collaborative task solving and simulating interaction to study emergent social behavior. We find that works mutually benefit from results in other categories: retrieval enables tool use, reflection improves multi-agent collaboration, and reasoning benefits all categories.
Conclusions: We discuss applications of agentic LLMs and provide an agenda for further research. Important applications are in medical diagnosis, logistics and financial market analysis. Meanwhile, self-reflective agents playing roles and interacting with one another augment the process of scientific research itself. Further, agentic LLMs provide a solution for the problem of LLMs running out of training data: inference-time behavior generates new training states, such that LLMs can keep learning without needing ever larger datasets. We note that there is risk associated with LLM assistants taking action in the real world—safety, liability and security are open problems—while agentic LLMs are also likely to benefit society.
@@ Abstract.
- 6 Reproducibility Checklist for JAIR
1. Introduction
The strength of the language abilities of LLMs has taken the world by storm. Recent work has extended their abilities with reasoning, information retrieval, and interaction tools. As a result, LLMs are now increasingly able to act as agents in the world (shen2024llm; qin2023toolllm). This ability has increased the relevance of LLMs to society and science. Agentic LLMs are being used to assist in medicine, logistics, finance, and other application areas. Their ability to self-reflect, interact, and play roles enables new types of research, including large-scale social science simulations. We survey the growing body of literature on agentic LLMs, which we define as large language models that (1) reason, (2) act, and (3) interact. We organize this article accordingly.
Agentic LLMs are also relevant in the acquisition of new training data for artificial intelligence (AI). Traditionally, LLMs have been trained on large datasets. Recently, however, it is getting harder to scale and improve datasets further, and training performance is reportedly plateauing, at high energy cost (sutskever2024). By interacting with the world, agents generate new empirical data (see Figure 1). This data can be used for additional training (pretraining or finetuning) or to enhance performance at inference time, provided there is adequate grounding through human or automated validation and filtering (subramaniam2025multiagent). An example of how LLMs can be trained based on their own actions, are Vision-Language-Action models, that update weights according to robotic action-feedback sequences (black2024pi_0; chiang2024mobility; yang2025magma). Thus, in addition to enabling useful applications, a second driver of interest in agentic LLMs is the opportunity to generate more training data.111Cognitive science teaches us that humans become more intelligent through interaction with the world and with other humans (we learn new behaviors and ideas from others) (brody1999intelligence; y2025intelligence). Societies of agents allow agentic LLMs to become more intelligent through interaction, as we will see in Section 4.
Agentic LLMs depend on progress in natural language processing, reasoning models, tool integration, reinforcement learning, agent-based modeling, and social science. At the confluence of these fields much exciting research has emerged.
This paper makes the following contributions:
@@ 1. Introduction
1.1. Agentic LLMs: Reasoning–Acting–Interacting
@@
Agency is a central concept in artificial intelligence (russell2016artificial). Agency is about identity and control, and about the capability to act on one’s goals or will (epstein1996growing; gilbert2019agent; barker2016cultural). Agents are endowed with decision-making capabilities, they sense changes in the environment, communicate, and act upon those changes (wooldridge1999intelligent), see also Figure 2. Agents have been studied for a long time, and occur in many fields of AI. From the definition of agents interacting with the environment, different approaches focus on specific aspects of agents and agent behavior. In symbolic reasoning (harman1984logic) and game theory (von2007theory; owen2013game), the topic of study is decision making by rational agents. The field of multi-agent systems studies intelligent systems that emerge from the interaction with different agents, human and/or artificial (ferber1999multi; steels2003intelligence). In machine learning, the field of reinforcement learning studies how an agent can learn from interacting with an environment (sutton2018reinforcement). In this context, agents are systems that would adapt their policy if their actions influenced the world in a different way (kenton2023discovering). In autonomous systems and robotics, agents act in order to achieve a goal (liu2018multiagent). Connectionism studies the emergence of intelligent behavior by embodied agents (brooks1990elephants; medler1998brief). Evolutionary algorithms (yu2010introduction; back1996evolutionary; preuss2015multimodal) use nature-inspired agent-based computation in order to achieve robust and flexible optimization, of which the ant-colony optimization algorithm (dorigo2007ant) is a well known example.
For the purpose of this survey, we build upon the definitions from these traditions. We define agentic LLMs as:
Agents that receive input in natural language from their environment, reason to make decisions, and take autonomous actions in affecting their environment, to achieve specific goals.
We stress that: agents may receive input and reason in natural or formal language; agents may plan to break down complex goals into smaller steps; agents may reflect on their own actions; agents may use tools to retrieve new information or to effect their actions; agent may build an internal model of the world; agents may have an internal structure that consists of multiple agents; agents may assist humans in achieving their goals; agents may interact in a society with humans and other agents; agents may create their own training data.
@@
The categories reasoning–acting–interacting build upon each other: the technology that has been developed by the reasoning approaches (category 1) is used for increasingly intelligent acting by assistants. The interactive abilities of the assistants (category 2) enable social simulation experiments. The outcome of assistant actions (category 2) and of these social experiments (category 3) can be used for data augmentation (category 1), to finetune LLMs (which can improve the accuracy of reasoning LLMs, etc.). This virtuous circle is depicted in Figure 1, and attracts interest from LLM researchers to agentic LLM methods (sutskever2024; guo2025deepseek; team2025kimi; lambert2024tulu). The categories also correspond to fields in artificial intelligence that have a long research tradition across symbolic AI, robotics/autonomous systems, and connectionism/multi-agent modeling, respectively. Agentic LLMs are thus both a recent development and build on decades of research. This is reflected in our discussion below.
1.2. Literature Selection
@@
Related surveys on agents and LLMs are starting to appear. li2024review reviews retrieval and tool use in agentic LLMs. wang2024survey focuses on autonomy and agent construction. gao2024large also provide an extensive overview, and focus on multi-agent modeling and simulation. xi2023rise again focus on the construction of interactive agents, using a more explanatory anthropomorphic approach of perception, brain, and action. An extensive general survey of LLMs is zhao2023survey, a slightly smaller one is (minaee2024large), an earlier survey is (min2023recent). yin2023survey review works on multimodal LLMs.
@@ 1.3. LLM Training Pipeline
Originally, language models used recurrent architectures such as LSTMs (hochreiter1997long) to embed semantic relations between token structures, allowing limited connections between tokens. The transformer architecture is an effective implementation of the attention mechanism (vaswani2017attention), allowing efficient random connections between tokens, improving performance greatly. Encoder transformer models, such as the BERT family (devlin2018bert), learn embeddings that are suitable for text understanding and classification. Decoder transformer models, such as the GPT family (brown2020language), are trained by masking for text completion and instruction following, and are suitable for text generation.
@@
LLMs are trained on large datasets (radford2019language; wei2022emergent). Performance on benchmarks testing formal linguistic competence is high (warstadt2019blimp) and so is accuracy on functional competence or natural language understanding tasks (GLUE, SQUAD, Xsum) (wang2018glue; wang2019superglue; rajpurkar2016squad; narayan2018don), translation (kocmi2022findings; papineni2002bleu; sennrich2015improving), and question answering (roberts2020much). Even in creative domains such as poetry and music composition LLMs have made some progress (zhang2024llm; yuan2024chatmusician; xing2025cocomposer).
@@
Popular LLMs are OpenAI’s ChatGPT series (achiam2023gpt; ouyang2022training), Meta’s LLaMa family (touvron2023llama), Anthropic’s Claude family (anthropic2024models), Google’s PaLM (chowdhery2023palm) and Gemini (team2023gemini), Qwen (yang2025qwen3), and the open-source models BLOOM (le2023bloom), Pythia (biderman2023pythia), OLMo (groeneveld2024olmoacceleratingsciencelanguage), and many others.
@@
LLMs are constructed using an elaborate pipeline with different training phases (radford2019language; minaee2024large). We will briefly describe the phases.
1. Acquire a large, general, unlabeled, text corpus (brown2020language).
2. Pretrain a transformer model on the corpus. This step yields a generalist natural language transformer model. The pretraining is done using a self-supervised attention approach (vaswani2017attention) on the unlabeled dataset (text corpus).
3. Finetune the general model to a specific (narrow) task using a supervised approach on a labeled dataset consisting of prompts and answers (supervised finetuning, SFT) (wei2022emergent; minaee2024large). This task can be, for example, translation from one language to another, or questions answering on a certain domain, such as medicine.
4. Instruction tune for improved instruction following. This is a form of supervised finetuning (ouyang2022training) to improve the ability to answer prompts.
5. Align the finetuned model with user expectations (preference alignment). The goal of this step is to improve the model to give socially acceptable answers such as prevention of hate speech. Popular methods are reinforcement learning with human feedback (RLHF) (ouyang2022training), direct preference optimization (DPO) (rafailov2024direct), or reinforcement learning with verifiable rewards (RLVR) (lambert2024tulu).
6. Optimize training to improve cost-effectiveness, for example with low-rank optimization (LoRa) (hu2021lora), mixed precision training (micikevicius2017mixed), or knowledge distillation (xu2024survey; gu2023minillm).
7. Infer using natural language prompts (instructions). This phase, inference, is the phase where, finally, we can use the fruits of our training efforts. Prompting is the preferred way of using LLMs. In LLMs whose size is beyond hundreds of billions parameters a new learning method emerges: in-context learning (brown2020language; wei2022emergent). This method provides a prompt that contains a small number of examples together with an instruction; it is a form of few-shot learning. However, no parameters of the model are changed by in-context learning, in-context learning takes place at inference time (dong2022survey; brown2020language).
Note that this pipeline is an example of a typical approach. Current pipelines are elaborate, and training is costly. Innovations to training pipelines are the topic of current research, see, for example, guo2025deepseek; team2025kimi; lambert2024tulu.
1.4. The Need for Agentic LLMs
@@
1. Prompt engineering Originally LLMs were trained as straight decoders, to be used with instruction prompts. The prompts contain context and instructions, and the model replies. The user interacts directly with the model, and writes the prompts themselves. LLMs turned out to be quite sensitive to small differences in the prompt formulation. When an answer is not satisfactory, the user has to remember the history of the interaction, and has to improve the prompt. This is known as prompt engineering. With basic LLMs, prompt improvement is a tedious, manual, task.
2. Hallucination When LLMs provide answers that look good, but are factually incorrect, they are said to hallucinate. Hallucination is a major problem of LLMs. It is caused, in part, by a lack of grounding. LLMs are trained to predict one of the statistically most probable next tokens, based on the training corpus. Since models are aligned to human preferences during fine-tuning, they often provide answers that look good by these standards while not adhering to other criteria, such as factuality. Various methods have been developed to mitigate hallucination, such as detecting uncertainty through self-reflection on their own answers, and with mechanistic interpretability methods (conmy2023towards). We will review papers that discuss these method in this survey.
3. Reasoning Another well-reported challenge for LLMs is (mathematical) reasoning (cobbe2021training; plaat2024reasoning). LLMs used to be quite bad at solving math word problems (such as: “Annie has a one pie that she cuts into twelve pieces. She eats one third of the pieces. How many pieces does she have left?”). Reasoning challenges have given rise to step-by-step problem solving methods, such as reported by wei2022chain, both implicit, and with explicit (neurosymbolic) prompt optimization methods (yao2024tree). This too we discuss in the next section.
4. Training Data LLMs are as smart as the data allows that was available at training time. When datasets no longer improve, pretraining and finetuning can no longer improve language models, and other learning methods are needed (sutskever2024). Any event that happened after training, or any information available in special databases, are not in the model (lewis2020retrieval).
@@ 1.5. Taxonomy
As we noted before, the three categories in our taxonomy come from three different backgrounds. To be intelligent, LLMs are enhanced with reasoning, combining deep learning with the symbolic AI tradition (yu2024distilling; li2025system). To be active, LLMs are enhanced with tools that can act in the world (including robots, that plan to move in the world). To be social, LLMs are placed in interactive settings with other agents. They rely partially on capacities already present in traditional LLMs, such as basic theory of mind abilities and understanding of game theory and social dilemmas. Agentic LLMs learn to interact better by adapting their intelligence.
@@
Earlier progress in multi-step reasoning LLMs and retrieval augmentation has enabled much of the current developments in agentic LLMs (category 1). In this category, the aim is to address challenges in solving math word problems, and in providing up to date answers to queries. The contributions to intelligent LLMs came out of the need to improve multi-step reasoning of basic LLMs (sub-category a). In reasoning LLMs, methods from planning and search are used to let the LLM follow a step-by-step reasoning path. More elaborate search algorithms allow automated prompt-improvement and self reflection (sub-category b). Finally, certain questions can only be answered by inference-time data retrieval (sub-category c). The focus is on the individual improvement of the intelligent LLM agent.
@@
Category 3 is about interaction in multi-agent simulations. Here, in sub-category a, we first study basic social capabilities of LLMs on which interactions can build. Second, in sub-category b, we study how LLM agents can work together using simulations where they are assigned specific roles. Third, in sub-category c, we study emergence of collective phenomena in open-ended interactions, such as social coordination via conventions and norms. Here the focus is on the emergent interaction level of the agent society. Multi-agent simulation with LLMs is becoming an active field for studying questions from the social sciences that previous generations of agent-based models were unable to address.
@@
| Chain of Thought (wei2022chain) | Step-by-step prompts | Math word (cobbe2021training) | Benchmark |
| Zero-shot CoT (kojima2022large) | ”Let’s think step-by-step” | Math word (cobbe2021training) | Benchmark |
| Self Consistency (wang2022self) | Ensemble | Math word (cobbe2021training) | Benchmark |
| Tree of Thoughts (yao2024tree) | depth-first-search prompts | Game of 24 | Benchmark |
| Implicit Planning (schultz2024mastering) | Train SoS (gandhi2024stream) | Chess, Hex | Benchmark |
| Progress Hint Prompt (zheng2023progressive) | Self Reflection | Math word (cobbe2021training) | Benchmark |
| Self Refine (madaan2023self) | Self Reflection | Dialogue Response | Benchmark |
| ReAct (yao2022react) | Reinforcement Learning | Decision Making | Benchmark |
| Reflexion (shinn2024reflexion) | Self Reflection/Reinf Learning | Decision Making | Benchmark |
| Self Discover (zhou2024self) | Self Reflection | Big Bench H (suzgun2022challenging) | Benchmark |
| Buffer of Thoughts (yang2024buffer) | Self Reflection | Math word (cobbe2021training) | Benchmark |
| Memory Coordination (zhang2023memory) | Self Reflection | LLM Personalization | Benchmark |
| Adaptive Retrieval (asai2023self) | Adaptive Retrieval | Question Answering | Benchmark |
| Retrieval Augmentation (lewis2020retrieval) | Retrieval Augmentation | Question Answering | Benchmark |
| MathPrompter (imani2023mathprompter) | Python Interpreter | Math problems | Benchmark |
| Program Aided Lang (gao2023pal) | Python Interpreter | Math word problems | Benchmark |
| Self Debugging (chen2023teaching) | Debugger | Code generation | Benchmark |
| FunSearch (romera2024mathematical) | Genetic Algorithm | Algorithm Generation | Benchmark |
| Planning Language (bohnet2024exploring) | Planner/PDDL | Blocksworld | Benchmark |
| Self taught Reasoner (zelikman2022star) | Reason augm finetuning | Math Word | Benchmark |
| DeepSeek R1 (guo2025deepseek) | Intrinsic Reasoning | Math Word | Benchmark |
@@
| WorldGPT (ge2024worldgpt) | Multimodal | World Model | WorldNet real-life scenarios |
@@
| WorldCoder (tang2024worldcoder) | Code model | World Code Model | Sokoban, MiniGrid, AlfWorld |
@@
| Task-planning (guan2023leveraging) | PDDL World model | Task finetuning | AlfWorld |
| CLIP (radford2021learning) | Multimodal | Vision Language | Benchmark |
@@
| Embodied BERT (suglia2021embodied) | Multimodal | Vision Language | ALFRED (shridhar2020alfworld) |
@@
| E2WM (xiang2024language) | Embodied World Model | MCTS + World Model | Question Answering |
| RT-2 (brohan2023rt) | Vision Language Action | VLA | Embodied reasoning tasks |
| LM-nav (shah2023lm) | Action traces | VLA | Topological navigation |
| Mobility VLA (chiang2024mobility) | long context demonstration | VLA | Navigation MINT |
| (black2024pi_0) | Flow Matching | VLA | Laundry folding, Table cleaning |
| Say Can (ahn2022can) | Grounded Actions | Value function for LLM | Manipulation, Kitchen |
@@
| Inner Monologue (huang2022inner) | Grounded Actions | Affordance in prompt | Manipulation, Kitchen |
@@
| Lang Guided Expl (dorbala2023can) | Generic Class Labels | Vision/Language | L-ZSON |
| Automatic Tool Chain (shi2024chain) | grounded reasoning | Tool behavior | ToolFlow |